OpenStudy (anonymous):

whats the numertator of the following rational expression e/f + 8/f = ?/f whats the numertator of the following rational expression e/f + 8/f = ?/f @Mathematics

OpenStudy (anonymous):

e+8

OpenStudy (anonymous):

You can multiply f on both sides, which gives you an equation of just numerators.

OpenStudy (anonymous):

e+8

OpenStudy (anonymous):

whats the numertator of the following rational expression p/q + 7/q = ?/q

OpenStudy (anonymous):

p+7

OpenStudy (anonymous):

whats the numertator of the following rational expression 3/k + 7/k = ?/k

OpenStudy (anonymous):

im sure u get the drift by now...

OpenStudy (anonymous):

is it k+7?

OpenStudy (anonymous):

no numerator means the top of the fraction. think again

OpenStudy (anonymous):

k+3

OpenStudy (anonymous):

3+7 = 10

OpenStudy (anonymous):

lmaoo ooo okay

OpenStudy (anonymous):

whats the numertator of the following rational expression a/b + 6/b = ?/b